### v1

# RadZero: Similarity-Based Cross-Attention for Explainable Vision-Language Alignment in Radiology with Zero-Shot Multi-Task Capability

Abstract
Recent advancements in multi-modal models have significantly improved vision-language alignment in radiology. However, existing approaches struggle to effectively utilize complex radiology reports for learning, rely on low-resolution images, and offer limited interpretability in attention mechanisms. To address these challenges, we introduce RadZero, a novel similarity-based cross-attention framework for vision-language alignment in radiology with zero-shot multi-task capability. RadZero leverages large language models to extract minimal semantic sentences from radiology reports and employs a multi-positive contrastive learning strategy to effectively capture relationships between images and multiple relevant textual descriptions. It also utilizes a pre-trained vision encoder with additional trainable Transformer layers, allowing efficient high-resolution image processing. By computing similarity between text embeddings and local image patch features, RadZero enables zero-shot inference with similarity probability for classification and pixel-level cross-modal similarity maps for grounding and segmentation. Experimental results on public chest radiograph benchmarks show that RadZero outperforms state-of-the-art methods in zero-shot classification, grounding, and segmentation. Furthermore, cross-modal similarity map analysis highlights its potential for improving explainability in vision-language alignment. Additionally, qualitative evaluation demonstrates RadZero’s capability for open-vocabulary semantic segmentation, further validating its effectiveness in medical imaging.
1 Introduction
Figure 1: Summary of RadZero’s capabilities. (a) Zero-shot multi-task performance (b) Open-vocabulary semantic segmentation. The original CXR images and the meanings of the abbreviations can be found in the Appendix.
Recent advances in deep learning have significantly transformed medical imaging, leading to numerous studies proposing computer-aided diagnosis [15, 8, 23]. However, obtaining high-quality manual annotations for medical images remains a major challenge. In the natural image domain, the emergence of vision-language models [27, 37, 36] has significantly reduced the reliance on exhaustive manual labels. These models learn the relationships between language and vision using only image-text pairs without label supervision, demonstrating strong performance even on zero-shot tasks such as classification and retrieval. Following these advancements, vision-language modeling has been rapidly progressing in the field of medical imaging, including chest X-rays (CXRs). Several studies focused on learning effective representations [34, 39, 40], demonstrating zero-shot capabilities [19, 22] without the need for task-specific annotations.
Despite these promising advances, medical vision-language alignment still faces several challenges. Effectively utilizing complex radiology reports for learning remains difficult. Previous attempts to address the issue include local alignment at the word or token level [12, 1], extracting clinical entities [39, 34], and LLM-based prompt alignment [19]. However, these approaches still suffer from limitations, such as failing to segment text embeddings into appropriate semantic units or the need for random selection during training, which leads to inefficient learning. Additionally, prior studies focusing on zero-shot tasks [19, 34] have relied on low-resolution images, which degrade performance on fine-grained tasks requiring precise localization. An even more critical challenge is ensuring explainability, as medical imaging models require interpretable outputs for clinical use. While attention maps [34, 19, 39] and dot-product similarities [12] are commonly used explainable features, they are not suitable for image-text similarity measures, limiting their effectiveness in interpretability.
To address these challenges, we propose RadZero, a novel vision-language alignment framework in radiology with zero-shot multi-task capability. We employ multi-positive contrastive learning [20] to effectively leverage multiple sentences associated with each image in an image-report pair. To use high-resolution images for improved performance on fine-grained zero-shot tasks, we adopt the idea of freezing a pretrained image encoder [36] and adding trainable Transformer layers for efficient training [18]. The key innovation of RadZero is the similarity-based cross-attention mechanism, which directly computes the similarity between text descriptions and local image patches. Cross-modal similarity map analysis demonstrates that these maps significantly enhance explainability by providing clear and interpretable visual reasoning for the model’s decisions. Additionally, we demonstrate the potential for open-vocabulary semantic segmentation, which can be achieved by simply applying a threshold to the similarity map. Experimental results on public chest radiograph benchmarks show that RadZero outperforms state-of-the-art (SOTA) models in zero-shot classification, grounding, and segmentation. Figure 1 summarizes the capabilities demonstrated by RadZero.
2 Related Works
2.1 General vision-language alignment
Contrastive learning for vision–language alignment with large-scale image–text pairs has been actively studied. CLIP [27] demonstrated that this approach enables strong zero-shot classification by directly aligning images and text. LiT [36] proposed freezing the pretrained vision encoder during contrastive training, preserving fine-grained visual features and further improving zero-shot performance. dino.txt [18] extended this paradigm by incorporating additional Transformer [31] layers on top of a pretrained DINOv2 [25], training only a lightweight module while keeping the vision encoder frozen. Additionally, it fused global and patch-averaged embeddings, enabling patch-level similarity computation with text and facilitating open-vocabulary semantic segmentation. UniCLIP [20] introduced a multi-positive NCE (MP-NCE) loss, which independently computes the contribution of multiple positive pairs per image. Building on these advances, our approach integrates a frozen, fine-grained vision encoder with trainable Transformer layers, following strategies from LiT and dino.txt. Additionally, we leverage MP-NCE loss to effectively align images with multiple text representations.
Figure 2: The overall framework of RadZero. (a) Finding-sentence extraction using an LLM. (b) Computation of the similarity logit, $l_{ij}^{n}$, between image $I_{i}$ and finding-sentence $S_{j}^{n}$. W-sum and cos-sim denote weighted sum and cosine similarity, respectively. (c) Computation of MP-NCE loss ($\mathcal{L}_{I}$) and InfoNCE loss ($\mathcal{L}_{T}$) from the similarity logit matrix.
2.2 Vision-language alignment in radiology
ConVIRT [40] initially applied contrastive learning to align CXR images with radiology reports. GLoRIA [12] focused on local alignment by introducing cross-attention between word-level text embeddings and patch-level image features, whereas MGCA [32] proposed a multi-granularity approach to capture relationships at the disease, instance, and pathological region levels. BioViL [3] refined the language model architecture to better handle radiology reports and BioViL-T [1] extended further by incorporating prior images during training. MedKLIP [34] utilized RadGraph [14] to extract triplets from reports and incorporated them into the training process, achieving notable performance in both zero-shot and fine-tuning settings. Similarly, KAD [39] employed RadGraph for entity extraction and used contrastive loss along with disease-specific queries in a cross-attention framework. More recently, CARZero [19] introduced an LLM-based prompt alignment to standardize diagnostic expressions in radiology reports, leveraging cross-attention alignment to achieve reliable zero-shot classification and grounding performance.
2.3 Fine-grained similarity for explainability
Although explainability is essential in medical vision-language alignment, the explainable features for image-text similarity proposed by previous studies have been insufficient. BioViL [3] computed patch-level similarity but without cross-attention on visual patches, limiting spatial alignment. GLoRIA [12] and BioViL-T [1] relied on word- or token-level embeddings, which lack semantic richness. Attention maps, though selected as explainable features by numerous zero-shot studies [39, 34, 19], do not constitute a valid measure of similarity between image and text. Due to the softmax activation, even unrelated image–text pairs yield high values at certain points in the map. The simple removal of softmax to use the raw logits is not a desirable solution, as the logits are neither scaled nor centered at zero. Moreover, the vector norms of query and key embeddings vary according to the image or text, causing fluctuations in the similarity value scale across different image–text pairs, an issue that dot-product similarity-based approaches [12] also encounter.
In contrast, our approach significantly improves explainability by employing similarity-based cross-attention between visual patches and text embeddings extracted from minimal semantic units via an LLM. The resulting feature maps are clear measures of fine-grained text–image similarity, providing intuitive and consistent explanations.
3 Methods
3.1 Finding-sentence extraction
Radiology reports contain a mixture of various types of information, including clinical history, observations, comparative analysis with prior studies, and diagnostic impressions. When using image and full report pairs for training, it becomes challenging to encode the complex radiology report into a single text embedding. CARZero [19] addressed this issue by leveraging an LLM to extract sentences from radiology reports and proposed a prompt alignment strategy using the template “There is [disease]” to ensure consistency between training and inference. Similarly, we utilize an LLM to extract such sentences, which we refer to as finding-sentences. We design a prompt that includes the positional information of the findings within a predefined template, such as “There is [finding] of [location]”. Each finding-sentence is segmented into the minimal semantic unit that contains the name of the finding, its presence (or uncertainty), and location information. The prompt used for finding-sentence extraction is in Appendix D. For each image, multiple finding-sentences are matched and used during training. An example of the finding-sentence extraction is illustrated in Figure 2 (a). For zero-shot inference, we apply prompt alignment by prepending “There is” to text descriptions such as findings and anatomical regions.
3.2 Vision-language alignment with similarity based cross-attention
3.2.1 Model architecture
To leverage the advantages of vision encoder pre-training, we adopt the approach of LiT [36] by freezing a pre-trained vision encoder in contrastive learning. In Vision Transformers [9] such as DINOv2 [25], interpolating the positional embeddings allows for increased input image resolution [29]. Building on this property, we train our model with high-resolution images. To embed the output of the vision encoder, we add trainable Transformer layers, as proposed by Jose et al. [18]. For the text encoder, we use a pre-trained Sentence-BERT [28], which is fine-tuned during training, to extract embeddings for each finding-sentence. The model architecture is illustrated in Figure 2 (b).
Figure 3: Numpy-like pseudocode implementation of similarity-based cross-attention.
3.2.2 Similarity-based cross-attention
To address the issues discussed in Sec. 2.3, we propose a cosine similarity-based cross-attention mechanism for computing similarity logits. By directly employing cosine similarity between the text and visual patch embeddings, we obtain cross-modal similarity scores that are well-defined in range and centered at zero. This consistent scaling allows for fair comparisons across different image–text pairs and significantly enhances explainability through the visualization of similarity maps. It also enables single thresholding, new possibilities for open-vocabulary semantic segmentation.
Specifically, let the size of a mini-batch be $B$. The $i$-th image $I_{i}$ ($i\in\{1,\dots,B\}$) is paired with $N_{i}$ finding-sentences. For $I_{i}$ and the $n$-th finding-sentence of the $j$-th image $I_{j}$ ($n\in\{1,\dots,N_{j}\}$), denoted by $S_{j}^{n}$, the algorithm to compute the patch-level similarity map $M_{ij}^{n}$ and the similarity logit $l_{ij}^{n}$ is summarized in Figure 3 as a numpy-like pseudocode.
From the vision encoder, we obtain a vision token sequence $V_{i}$, composed of the [CLS] token embedding and visual patch embeddings. The subsequent cross-attention operation takes $T_{j}^{n}$, the text embedding of $S_{j}^{n}$, as the query and $V_{i}$ as both the key and value. We compute the cosine similarity between $T_{j}^{n}$ and each token in $V_{i}$, and to ensure distinguishability, we multiply it by a temperature parameter. The softmax activation transforms these scaled similarity scores into attention probabilities. The weighted sum of the attention probabilities and the vision tokens yields a refined vision embedding. Finally, we compute its cosine similarity with $T_{j}^{n}$, multiply it by a temperature parameter, and obtain the overall similarity logit $l_{ij}^{n}$. Additionally, we obtain the patch-level cross-modal similarity map $M_{ij}^{n}$ by excluding the first element (corresponding to the [CLS] token) from the similarity scores.
3.3 Training objectives
Although CARZero also uses prompt templates for training, it suffers from instability due to randomly selecting one sentence for each image at every training step. To utilize all $N$ finding-sentences matched to each image at every step, we adopt multi-positive NCE (MP-NCE) loss [20] which treats positive pairs independently in order to amplify the loss contributions from each positive pair. A visualization of our contrastive loss can be seen in Figure 2 (c). Let $N_{T}=\sum_{i=1}^{B}N_{i}$ be the total number of finding-sentences in a mini-batch. For the $i$-th image, the number of positive and negative finding-sentences are $N_{i}$ and $N_{T}-N_{i}$, respectively. The MP-NCE loss can be computed as follows:
$\mathcal{L}_{I}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{e^{l_{ii}^{n}}}{e^{l_{ii}^{n}}+\sum\limits_{j\neq i}^{B}\sum\limits_{m=1}^{N_{j}}e^{l_{ij}^{m}}}$ (1)
For each finding-sentence $S_{i}^{n}$, there is one positive image $I_{i}$ and $B-1$ negative images. The corresponding InfoNCE loss [24] is computed as follows:
$\mathcal{L}_{T}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{e^{l_{ii}^{n}}}{e^{l_{ii}^{n}}+\sum\limits_{j\neq i}^{B}e^{l_{ji}^{n}}}$ (2)
The final objective function is the sum of $\mathcal{L}_{I}$ and $\mathcal{L}_{T}$:
$\mathcal{L}=\mathcal{L}_{I}+\mathcal{L}_{T}.$ (3)
Figure 4: Zero-shot inference pipeline of RadZero.
3.4 Zero-shot inference
The similarity logit between an image $I_{i}$ and a text sentence $S_{j}^{n}$, denoted by $l_{ij}^{n}$, is converted into the final similarity probability $\hat{l}_{ij}^{n}$ by applying a sigmoid function, which is then used for classification. For zero-shot grounding and segmentation, pixel-level probability values are required. Since the patch-level cross-modal similarity map $M_{ij}^{n}$ is computed after image pre-processing such as padding and resizing, we restore it to the original size using linear interpolation while accounting for these transformations. Then, a sigmoid function is applied to obtain the pixel-level cross-modal similarity map ${\hat{M}}_{ij}^{n}$, which we simply refer to as the similarity map. The zero-shot inference process is illustrated in Figure 4.
3.5 Explainability
In the cross-attention and logit computation process, we do not modify the embedding space beyond applying $L_{2}$ normalization and adjusting the temperature. Since similarity is computed using cosine similarity between vision patches and text embeddings, the final similarity map ${\hat{M}}_{ij}^{n}$ can be directly interpreted as the similarity between each image pixel and the text, significantly improving the model’s explainability. For instance, when the similarity between a text prompt and an image is low, it directly reflects weak alignment between the vision token sequence and the text embedding, as the final similarity logit is derived from a weighted sum of vision tokens. Conversely, when the similarity is high, it indicates that the text embedding closely aligns with either the entire vision token sequence or specific visual patches. This approach allows for a transparent visualization of how similarity logits are computed through the similarity map. In the context of disease diagnosis, the model can explicitly reveal how conclusions are derived, greatly enhancing its explainability and making its decision-making process more interpretable.
4 Experiments
4.1 Training dataset
MIMIC-CXR [16]
We train our model using the MIMIC-CXR dataset for vision-language alignment. MIMIC-CXR comprises 377,110 CXR images from 227,835 radiographic studies involving 65,379 patients. Each study includes a radiology report and one or more CXR images in either frontal or lateral views. Images are sourced from MIMIC-CXR-JPG [17], and only the findings and impression sections of reports are extracted using the official codebase [^1] https://github.com/MIT-LCP/mimic-cxr. All view positions are considered, and the official dataset split is followed. As described in Section 3.1, finding-sentence extraction is applied, with each study containing an average of 6.45 such sentences. Studies without extracted finding-sentences are discarded, resulting in 352,875 training images and 2,852 for validation.
4.2 Test datasets
Open-I [7] dataset consists of 3,851 radiology reports and 7,470 CXR images, with multi-label classification annotations for 18 diseases. PadChest [4] consists of 160,868 CXR images collected from 67,000 patients and exhibits a long-tailed distribution with a total of 192 labels. Following [19], we use 39,053 samples annotated by board-certified radiologists. Additionally, PadChest20, introduced in [19], serves as a test set for rare disease evaluation, consisting of 20 classes with fewer than 10 samples each. ChestXray14 [33] provide official test set with 22,433 images and corresponding labels for 14 diseases. CheXpert [13] provide official official test set consists of 500 patients’ images annotated by five board-certified radiologists. Following [19], we perform classification evaluation on five observations: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. ChestXDet10 [21] is a subset of ChestXray14, consisting of 542 images with bounding box annotations for 10 diseases in the official test set. SIIM [35] pneumothorax dataset contains manually annotated segmentation masks for CXRs, of which 11,582 images are publicly available. We adopt the test split of [32], comprising 1,704 images with 458 positive samples. RSNA [30] pneumonia dataset consists of 29,700 frontal-view radiographs with bounding box annotations indicating evidence of pneumonia. We use the test set of 5,337 images released by [34], of which 1,218 are positive. MS-CXR [2] consists of 1,153 image-phrase-bounding box triplets, with images sourced from MIMIC-CXR. The bounding boxes annotated to specific phrases in the report enable more detailed grounding, referred to as phrase grounding. For a fair comparison on the test set of 167 images released by [6], where each phrase corresponds to only one bounding box, we exclude these test images from the training set described in Sec. 4.1.
4.3 Evaluation metrics
AUC, or area under the ROC curve, is adopted to evaluate zero-shot classification on multi-label test datasets. Pointing game [38], which determines whether the coordinates of the maximum value falls within the corresponding bounding box, is employed as the grounding metric. Dice score serves as a standard evaluation metric for segmentation. Following [34], we compute the Dice score using only positive samples and optimize the segmentation threshold to maximize the score. Threshold search intervals are 0.01 for sigmoid and 0.001 for softmax, depending on the feature map’s activation function. Pixel-wise AUC (Pix-AUC) computes AUC at pixel-level to evaluate the quality of the segmentation probability map. To account for both sensitivity and specificity in mask prediction, we incorporate both positive and negative samples. For fine-grained tasks such as grounding and segmentation, predictions are interpolated back to the original image size before evaluation.
4.4 Implementation details
We adopt XrayDINOv2 [5] as the pre-trained vision encoder, which is trained in a unimodal setting using CXR images based on DINOv2 [25]. While the vision encoder was trained with an image resolution of 224, we increase it to 518 for our experiments. The patch size of $14\times 14$ leads to $37\times 37$ patches, yielding a vision patch length $L$ of 1369. The text encoder is MPNet (“all-mpnet-base-v2”) [28], initialized with pre-trained parameters. The Transformer layers are randomly initialized, with a hidden dimension of 768, matching that of both the vision and text encoders. While the vision encoder remains frozen, all other parameters are trainable. Following [27], the learnable temperature parameter is set to $\log(1/0.07)$. The details of model training can be found in Appendix B. The LLM used for extracting finding-sentences is “Llama-3.3-70B-Instruct” [10].
5 Results
In Sec. 5.1, we compare RadZero with other SOTA approaches in zero-shot classification, grounding, and segmentation tasks. In Sec. 5.2, we analyze the model’s explainability through cross-modal similarity map analysis. In Sec. 5.3, we validate the potential of RadZero for open-vocabulary semantic segmentation. Due to space limitations, the full ablation study is provided in Appendix C.
  Method Open-I PadChest PadChest20 ChestXray14 CheXpert ChestXDet10 SIIM RSNA
GLoRIA [12] 0.589 0.565 0.558 0.610 0.750 0.645 - -
BioViL-T [1] 0.702 0.655 0.608 0.729 0.789 0.708 - -
MedKLIP [34] 0.759 0.629 0.688 0.726 0.879 0.713 0.897 0.869
KAD [39] 0.807 0.750 0.735 0.789 0.905 0.735 - -
CARZero [19] 0.838 0.810 0.837 0.811 0.923 0.796 0.924 0.747
RadZero 0.847 0.841 0.871 0.804 0.900 0.787 0.924 0.834
Table 1: Comparison of zero-shot classification performance between RadZero and baseline models on various public CXR datasets. The metric is AUC score, with the best results highlighted in bold.
  Method Mean ATE CALC CONS EFF EMPH FIB FX MASS NOD PTX
GLoRIA [12] 0.367 0.479 0.053 0.737 0.528 0.667 0.366 0.013 0.533 0.156 0.143
KAD [39] 0.391 0.646 0.132 0.699 0.618 0.644 0.244 0.199 0.267 0.316 0.143
BioViL-T [1] 0.351 0.438 0.000 0.630 0.504 0.846 0.390 0.026 0.500 0.000 0.171
MedKLIP [34] 0.481 0.625 0.132 0.837 0.675 0.734 0.305 0.224 0.733 0.312 0.229
CARZero [19] 0.543 0.604 0.184 0.824 0.782 0.846 0.561 0.184 0.700 0.286 0.457
RadZero 0.622 0.646 0.368 0.824 0.857 0.872 0.585 0.250 0.767 0.506 0.543
Table 2: Comparison of zero-shot grounding performance between RadZero and baseline models on ChestXDet10. The metric is pointing game accuracy, with the best results highlighted in bold. Lesion abbreviations can be found in the Appendix A.
  Method MS-CXR
BioViL-T [1] 0.719
MedKLIP [34] 0.407
CARZero [19] 0.749
RadZero 0.844
Table 3: Comparison of zero-shot phrase grounding performance between RadZero and baseline models on MS-CXR. The metric is pointing game accuracy.
  Method RSNA SIIM
Dice Dice Pix-AUC
GLoRIA [12] 0.347* - -
BioViL [3] 0.439* - -
MedKLIP [34] 0.465* 0.044 0.648
CARZero [19] 0.540 0.100 0.856
CARZero (logits) 0.529 0.081 0.928
RadZero 0.546 0.171 0.947
  MGCA [32] (1%) 0.513 0.144 0.752
MGCA (10%) 0.571 0.238 0.856
MGCA (100%) 0.578 0.305 0.976
Table 4: Comparison of zero-shot segmentation performance between RadZero and baseline models on RSNA and SIIM. Values marked with * are from Wu et al. [34], while we conduct evaluations for other scores. The best results in zero-shot setting are highlighted in bold.
5.1 Zero-shot evaluation
Classification.
Table 1 compares RadZero with other SOTA models in public test datasets. For the five datasets reported by CARZero [19], we extracted the reported results, while for SIIM and RSNA we independently evaluated two open-source models. Notably, RadZero achieved new SOTA performance on Open-I and PadChest. In PadChest, which is a long-tailed dataset containing 192 classes, we outperformed CARZero by 3.1 points, demonstrating a strong generalization in zero-shot classification. We also observe notable gains in PadChest20 (a subset focusing on rare diseases), suggesting that similarity-based vision-language alignment is especially effective in handling less frequent conditions. For the datasets in which our model missed the first place, MedKLIP performed the best on RSNA, while CARZero excelled on ChestXray14, CheXpert, and ChestXDet10. However, MedKLIP underperformed on the latter datasets, and CARZero underperformed on RSNA. In contrast, RadZero showed results comparable to the top-performing models across all four datasets.
The representative classification metric shown in Figure 1 is the average AUC across all datasets. RadZero established a new SOTA, outperforming CARZero by 1.5 percentage points. This improvement stems from our proposed training strategy, which leverages finding-sentences as minimal semantic units in contrastive learning and incorporates multi-positive training to enhance the diversity of both positive and negative samples per image.
Grounding.
Table 2 presents the results of zero-shot grounding on ChestXDet10. We adopted the pointing game scores reported by CARZero for each disease, except for BioViL-T, which we evaluated using the released model. Based on the average scores across all diseases, RadZero achieved the highest score, surpassing CARZero by a margin of 0.079. Further analysis of the results for each lesion revealed that our model demonstrated the best performance across all classes, with the exception of consolidation. This indicates that our similarity map effectively captures the local similarity between text and visual patch embeddings, irrespective of the disease class. Additionally, our architecture can efficiently train with higher input resolutions, enabling more precise grounding.
Table 3 presents the results of zero-shot phrase grounding. The model’s ability to localize the image region corresponding to a text phrase was measured using pointing game accuracy. We evaluated all baselines using publicly available models. RadZero achieved the highest score of 0.844, showing that it accurately interpreted text phrases. This advancement can be attributed to the training process, which effectively learned the fine-grained similarity between image and text, and the utilization of location information during the finding-sentence extraction process.
Segmentation.
Table 4 summarizes the segmentation results on SIIM and RSNA. To compare zero-shot performance with supervised models, we fine-tuned the pre-trained MGCA model, with percentage values in parentheses indicating the proportion of training data used.
Among zero-shot models, RadZero achieved the best Dice scores on both SIIM and RSNA. Notably, on SIIM, it showed a 71% improvement of Dice score over CARZero, demonstrating superior segmentation capability. However, the gap in RSNA is smaller, which can be attributed to mismatch in granularity—RSNA’s labels are bounding boxes rather than pixel-wise annotations, which hinders the merit of RadZero’s fine-grained similarity prediction.
RadZero remains competitive even when compared to fine-tuned models. It outperformed MGCA (1%) on both SIIM and RSNA, demonstrating the potential of zero-shot segmentation. While its performance is lower than MGCA (10%) and MGCA (100%), RadZero does not require mask labels during training, allowing it to generalize beyond a fixed vocabulary. This advantage is further demonstrated in our open-vocabulary segmentation results in Sec. 5.3.
SIIM contains fine-grained mask labels, making it suitable for evaluating Pix-AUC scores. RadZero achieved the highest Pix-AUC score, surpassing even MGCA (10%), demonstrating its capability to generate well-calibrated similarity maps that effectively distinguish between positive and negative regions. In contrast, MedKLIP and CARZero, which rely on attention maps, resulted in lower Pix-AUC scores of 0.648 and 0.856, respectively. For a fair comparison, we also evaluated CARZero’s pre-softmax logit maps (CARZero (logits)), which improved results but still fell short of RadZero. CARZero (logits) underperformed even its own original Dice score (0.081 vs. 0.100) despite the optimal threshold search, which aligns with our discussion in Sec. 2.3: dot product-based similarity lacks consistent scaling. In contrast, RadZero’s mask prediction map directly represents text–image pixel-level similarity, allowing low similarity for negative samples. This is reflected in its high Pix-AUC score, further supported by our similarity map analysis in Sec. 5.2.
Figure 5: Visualization of cross-modal similarity maps The CXR images of (a), (b), and (c) are from ChestXDet10, representing normal, fibrosis, and effusion in the right lung, respectively. The similarity probability $\hat{l}$ (top-right corner) between each CXR image and the text prompt (bottom-right corner) can be explained through the visualized similarity map $\hat{M}$.
5.2 Cross-modal similarity map analysis
Figure 5 demonstrates that RadZero effectively aligned visual and textual representations through similarity-based cross-attention. Its outputs, the similarity map $\hat{M}$ and probability $\hat{l}$, offer interpretable visualizations as well as quantitative metrics. For the normal image (Figure 5 (a)), the model assigned low similarity (0.072) to the prompt “There is atelectasis,” with a dark similarity map, indicating no strong alignment. In contrast, the prompts “There is no atelectasis” (0.921) and “The lungs are clear” (0.846) produced high similarities, with bright activations in the lung fields, confirming alignment with normality.
Figure 5 (b) shows fibrosis, and “There is fibrosis” resulted in high similarity (0.970) with strong activations in the affected lung. Conversely, RadZero produced much lower scores (0.093 and 0.014) and darker similarity maps for the two prompts indicating normality, effectively distinguishing between normal and abnormal descriptions.
The key strength of our approach is its ability to differentiate anatomical descriptions. For the right-sided pleural effusion in Figure 5 (c), the model assigned high similarity (0.907) to “There is effusion,” with bright activations in the correct region. Notably, “There is right effusion” (0.946) scored even higher, indicating accurate localization, while “There is left effusion” (0.125) resulted in a much lower score and a dark similarity map, showing that the model correctly distinguishes between the left and right lungs.
Overall, these results highlight the explainability of RadZero. The similarity probability can be validated at the pixel-level, ensuring spatially grounded explanations. Thus, RadZero provides enhanced interpretability by offering a transparent rationale for how conclusions are derived.
5.3 Open-vocabulary semantic segmentation
Figure 6: Open-vocabulary semantic segmentation for findings. The CXR images and bounding box labels are sourced from ChestXDet10. The segmentation threshold was set to 0.7.
Figures 6 and 7 show the results of open-vocabulary semantic segmentation for findings and anatomical regions, respectively. Segmentation mask was obtained for each text prompt by thresholding the similarity map $\hat{M}$. When multiple prompts were identified as positive for a pixel, the prompt with the highest similarity was assigned to it.
Figure 6 shows that the model effectively identified lesion locations based on text prompts. Some regions extended beyond ground truth bounding boxes, clearly indicating room for improvement. Yet, among these incorrect predictions were a few instances where the model captured clinically relevant features that were not explicitly annotated. In Figure 6 (b), RadZero captured a chest tube as similar to “pneumothorax”, which is a reasonable association.
Figure 7 illustrates RadZero’s capability to segment anatomical regions without explicit supervision. Although the model struggles to delineate precise boundaries, it identifies approximate regions, suggesting the ability to infer spatial relationships from text. However, occasional misclassifications, such as identifying ribs as the spine, underscore its limitations and the need for further refinement.
These results demonstrate the potential of RadZero for zero-shot open-vocabulary semantic segmentation, aligning text descriptions with medical images. This approach offers a promising direction for interpretable medical image segmentation, with further advancements anticipated. Additional qualitative results are provided in Appendix E.
Figure 7: Open-vocabulary semantic segmentation for anatomical regions. The CXR images are sourced from Open-I. The segmentation threshold was set to 0.4.
6 Conclusion
In this work, we introduced RadZero, a novel similarity-based cross-attention framework for vision-language alignment in radiology. By computing cosine similarities between text descriptions and local image patches, RadZero enhanced interpretability while demonstrating remarkable zero-shot capability across classification, grounding, and segmentation. Its ability to process high-resolution chest X-rays combined with a multi-positive contrastive learning strategy, enabled effective representation learning without requiring pixel-level annotations. Extensive evaluations on public chest radiograph benchmarks revealed RadZero’s superiority over state-of-the-art methods in zero-shot classification, grounding, and segmentation. Furthermore, cross-modal similarity map analysis highlighted its advantage in explainability, as the similarity map provide a transparent rationale for how conclusions are derived. Qualitative assessments further demonstrated RadZero’s potential for open-vocabulary semantic segmentation, validating its adaptability to diverse medical imaging scenarios.
RadZero, while achieving impressive results, has limitations that indicate areas for future research. It did not consistently surpass all benchmarks in zero-shot classification, emphasizing the need for improved generalization. Fine-tuning with class labels could facilitate the development of a high-performance explainable classifier. Extending the similarity-based cross-attention approach to imaging modalities like CT or MRI offers potential for interpretable vision-language models in medical imaging.
Supplementary Material
Appendix A Abbreviation
  Abbreviation Description
ATE Atelectasis
CALC Calcification
CONS Consolidation
EFF Effusion
EMPH Emphysema
FIB Fibrosis
FX Fracture
MASS Mass
NOD Nodule
PTX Pneumothorax
UL Upper Lobe
ML Mid Lobe
LL Lower Lobe
Rclav Right Clavicle
Lclav Left Clavicle
RULZ Right Upper Lung Zone
RMLZ Right Mid Lung Zone
RLLZ Right Lower Lung Zone
LULZ Left Upper Lung Zone
LMLZ Left Mid Lung Zone
LLLZ Left Lower Lung Zone
RCPA Right Costophrenic Angle
LCPA Left Costophrenic Angle
HD Hemidiaphragm
RHD Right Hemidiaphragm
LHD Left Hemidiaphragm
Table 5: Abbreviations for lesions and anatomical regions.
Appendix B Model Training Details
RadZero is trained for 20 epochs with an early stopping patience of 5 epochs, selecting the best model based on validation loss. We employ the AdamW optimizer with a learning rate of 0.0001, following a cosine decay scheduler, with 50 warm-up steps, a weight decay of 0.05, and gradient clipping set to 1.0. Training is conducted with a global batch size of 256 using distributed data parallel (DDP) on four H100 GPUs for 16 hours.
Appendix C Ablation Studies
We conducted ablation studies to assess the impact of key design choices by selectively modifying parts of our approach. The zero-shot tasks for evaluation included classification (class.), grounding (ground.), phrase grounding (phrase.), and segmentation (seg.). Classification was tested on the PadChest dataset, known for its highly imbalanced (long-tailed) label distribution, with AUC as the evaluation metric. Grounding and phrase grounding were evaluated using the pointing game on the ChestXDet10 and MS-CXR test sets, respectively. Segmentation performance was measured by the Dice score on the SIIM dataset. For the ablation study, the default batch size and maximum number of epochs were set to 128 and 10, respectively.
Impact of multi-positive pairs for contrastive learning.
To apply contrastive learning, CARZero [19] randomly sampled one sentence from multiple available options. We compare this random selection strategy with our multi-positive NCE loss, which leverages all sentences associated with an image. As shown in Table 6, utilizing multiple sentences consistently outperformed random selection across all tasks, demonstrating the effectiveness of our approach.
  Method class. ground. phrase. seg.
Random Select 0.827 0.586 0.832 0.138
Multi-positive 0.841 0.622 0.844 0.171
Table 6: Impact of multi-positive pairs for contrastive learning.
Impact of view position.
Table 7 shows the performance variation based on the view position of CXR images. We compared two models: one trained exclusively on frontal view images from MIMIC-CXR and another trained on both frontal and lateral views. The model trained on all view positions consistently outperformed the frontal-only model, suggesting that it effectively learned to interpret lateral images, enhancing overall robustness.
  View Position class. ground. phrase. seg.
Frontal 0.831 0.604 0.838 0.161
All View 0.841 0.622 0.844 0.171
Table 7: Impact of view position.
Impact of image encoder and resolution.
Table 8 presents the impact of the image encoder and image resolution on model performance. We compared M3AE [11], RadDINO [26], and XrayDINOv2 [5] using image resolutions of 224 and 518. For M3AE, we used the pre-trained model released by Lai et al. [19], and all encoders were frozen during training.
DINOv2-based encoders significantly outperformed M3AE across all tasks, demonstrating the effectiveness of the DINOv2 pretraining strategy for chest X-ray representation learning. Comparing XrayDINOv2 at resolutions of 224 and 518, we observe that higher image resolution improves fine-grained tasks such as grounding and segmentation. RadDINO and XrayDINOv2 showed similar performance, suggesting that our approach is effectively applied to models trained with the DINOv2 strategy on chest X-ray images.
  Image Image class. ground. phrase. seg.
Encoder Resolution
M3AE [11] 224 0.703 0.284 0.647 0.027
RadDINO [26] 518 0.850 0.610 0.844 0.144
XrayDINOv2 [5] 224 0.841 0.548 0.832 0.118
XrayDINOv2 518 0.841 0.622 0.844 0.171
Table 8: Impact of image encoder and resolution.
Impact of trainable vision layer architecture.
Table 9 presents the impact of different trainable layers in the image encoder. The commonly used linear layer showed relatively lower performance across tasks. In contrast, two transformer layers achieved the best results in classification, grounding, and phrase grounding.
Based on this observation, RadZero was designed with two transformer layers added to the vision encoder. This improvement is likely due to the transformer’s ability to attend to all patch embeddings, capturing richer semantic information.
  Model class. ground. phrase. seg.
Linear 0.826 0.549 0.826 0.100
1 Transformer layer 0.835 0.585 0.832 0.158
2 Transformer layer 0.841 0.622 0.844 0.171
Table 9: Impact of trainable vision layer architecture.
Impact of text encoder.
Table 10 presents the performance of different text encoders used during training. We compared MPNet and BioBERT, where BioBERT was fine-tuned on clinical reports by CARZero [19].
While MPNet showed slightly lower performance in classification, it achieved notable improvements in phrase grounding and segmentation, demonstrating its effectiveness in tasks requiring fine-grained text-image alignment.
  Image class. ground. phrase. seg.
Encoder
BioBERT 0.842 0.582 0.832 0.127
MPNet 0.841 0.622 0.844 0.171
Table 10: Impact of text encoder.
Impact of batch size.
Table 11 presents the impact of batch size on model performance during training. To ensure a fair comparison, we maintained a consistent total number of training steps by adjusting the number of epochs: 5 for a batch size of 64, 10 for 128, and 20 for 256.
We observed that a batch size of 64 resulted in lower performance across all tasks. While the model trained with a batch size of 128 performed reasonably well, its zero-shot grounding performance was notably lower than that of the 256 batch size model. As a result, we selected 256 as the final batch size.
This trend aligns with the well-known impact of batch size in contrastive learning, where larger batch sizes generally improve representation learning by providing more diverse negative samples, leading to better alignment and discrimination.
  Batch size class. ground. phrase. seg.
64 0.835 0.583 0.826 0.165
128 0.840 0.594 0.850 0.177
256 0.841 0.622 0.844 0.171
Table 11: Impact of batch size.
Figure 8: Prompt design for extracting finding-sentences with LLM.
Appendix D Prompt for finding-sentence extraction.
As shown in Figure 8, the prompt instructs the LLM to extract clinically relevant minimal semantic units in the form of sentences from radiology reports. Finding-sentences are standardized through prompt alignment to follow a “There is” format, with a one-shot example enhancing extraction accuracy and guiding the model to identify both findings and their corresponding anatomical locations in a structured manner.
Appendix E Additional Visualization Results
Figure 9 presents cross-modal similarity maps for 10 different findings, following the pipeline in Sec. 5.2. The highest similarity regions align well with the bounding boxes, even for multiple or small lesions. The similarity probability was above 0.5 for all findings except calcification. While the model correctly localized calcifications, the activated regions appeared as small bright spots, leading to a lower similarity probability of 0.45 due to the weighted sum calculation. This highlights a limitation of RadZero, suggesting the need for further refinement in future work.
Figure 10 and Figure 11 present open-vocabulary semantic segmentation results, following Sec. 5.3. These figures include images from Figure 1 for qualitative analysis, along with additional examples not shown in the introduction.
Figure 10 depicts segmentation of anatomical regions, which, while not perfect, generally align with appropriate locations. Figure 11 presents examples demonstrating RadZero’s potential for open-vocabulary semantic segmentation, including additional lesion types such as mass, fibrosis, and calcification.
Figure 9: Cross-modal similarity maps for 10 findings. Visualization of similarity maps generated by RadZero on the ChestXDet10 dataset. Red boxes indicate ground truth bounding boxes. The similarity probability $\hat{l}$ is shown in the top-right corner of each map.
Figure 10: Open-vocabulary semantic segmentation for anatomical regions. The CXR images are sourced from Open-I. The segmentation threshold was set to 0.4.
Figure 11: Open-vocabulary semantic segmentation for findings. The CXR images and bounding box labels are sourced from ChestXDet10. The segmentation threshold was set to 0.7.