### v2

# RadZero: Similarity-Based Cross-Attention for Explainable Vision-Language Alignment in Radiology with Zero-Shot Multi-Task Capability

Abstract
Recent advancements in multi-modal models have significantly improved vision-language (VL) alignment in radiology. However, existing approaches struggle to effectively utilize complex radiology reports for learning and offer limited interpretability through attention probability visualizations. To address these challenges, we introduce RadZero, a novel framework for VL alignment in radiology with zero-shot multi-task capability. A key component of our approach is VL-CABS (Vision-Language Cross-Attention Based on Similarity), which aligns text embeddings with local image features for interpretable, fine-grained VL reasoning. RadZero leverages large language models to extract concise semantic sentences from radiology reports and employs multi-positive contrastive training to effectively capture relationships between images and multiple relevant textual descriptions. It uses a pre-trained vision encoder with additional trainable Transformer layers, allowing efficient high-resolution image processing. By computing similarity between text embeddings and local image patch features, VL-CABS enables zero-shot inference with similarity probability for classification, and pixel-level VL similarity maps for grounding and segmentation. Experimental results on public chest radiograph benchmarks show that RadZero outperforms state-of-the-art methods in zero-shot classification, grounding, and segmentation. Furthermore, VL similarity map analysis highlights the potential of VL-CABS for improving explainability in VL alignment. Additionally, qualitative evaluation demonstrates RadZero’s capability for open-vocabulary semantic segmentation, further validating its effectiveness in medical imaging.
1 Introduction
Recent advancements in deep learning have significantly impacted medical imaging, leading to numerous studies on computer-aided diagnosis [15, 8, 24, 35]. However, acquiring high-quality manual annotations remains a key challenge. In contrast, vision-language (VL) models (VLMs) in the natural image domain [28, 39, 38] have reduced reliance on manual labeling by learning from image-text pairs without explicit supervision, achieving strong zero-shot performance in tasks like classification and retrieval. Building on this progress, VLMs have been increasingly explored in medical imaging, including chest X-rays (CXRs). Several studies have demonstrated effective representation learning [36, 41, 42] and zero-shot capabilities [19, 23] without task-specific annotations.
Figure 1: Comparison of attention maps and the proposed VL similarity map for visualizing VL alignment. (a) While traditional attention maps inevitably exhibit high values at certain points due to softmax activation, the proposed VL similarity maps yield low values for unrelated image-text pair. (b) Their fixed scale, originating from cosine similarity, enables open-vocabulary semantic segmentation through simple thresholding.
Despite these advance, current medical VLMs underutilize the rich semantics of radiology image–report pairs. Prior methods—such as word-level alignment [12, 1], clinical entity extraction [41, 36], and using LLM prompts [19]—face limitations, struggle with poor text embedding segmentation and inefficient training due to random sampling. An effective solution should 1) decompose reports into semantically minimal, clinically meaningful sentences, 2) embed each sentence independently, and 3) learn from multiple sentence–image pairs per study to to fully leverage supervision.
Moreover, reliable explainability is critical for the clinical adoption of medical VLMs. Attention maps are used as explainable features by most of the recent research [36, 19, 41]. However, while an attention map (Figure 1(a), top) can indicate where the model is focusing, it does not provide an explanation for why it is attending to those regions. This limitation can be addressed by computing pixel-level image–text similarity, which enables more fine-grained and transparent explanations.
Figure 2: Zero-shot multi-task performance. Each score is averaged over multiple datasets per task.
To overcome these shortcomings, we propose RadZero, a novel VL alignment framework for radiology with zero-shot multi-task capabilities. RadZero employs multi-positive contrastive learning [20] to incorporate multiple sentences per image-report pair. RadZero’s core innovation, Vision-Language Cross-Attention Based on Similarity (VL-CABS), directly computes cosine similarity between text descriptions and local image patches. Unlike traditional attention maps, the resulting VL similarity maps offer clearer visual reasoning by maintaining low values for unrelated image-text pairs (Figure 1 (a), bottom). This enhances interpretability and enables open-vocabulary semantic segmentation via simple thresholding. RadZero also supports high-resolution inputs by freezing a pre-trained image encoder [38] and adding trainable Transformer layers [18], further boosting performance on fine-grained zero-shot tasks. Experiments on public chest radiograph benchmarks demonstrate that RadZero outperforms state-of-the-art (SOTA) models in various zero-shot tasks (Figure 2), while qualitative analyses reveal its enhanced explainability and potential for open-vocabulary semantic segmentation.
2 Related Works
2.1 General vision-language alignment
Contrastive learning for vision–language alignment with large-scale image–text pairs has been actively studied. CLIP [28] demonstrated that directly aligning images and text enables strong zero-shot classification. LiT [38] proposed freezing the pre-trained vision encoder during contrastive training, preserving fine-grained visual features and further enhancing zero-shot performance. dino.txt [18] extended this framework by adding Transformer [32] layers on top of a pre-trained DINOv2 [26], training only a lightweight module while keeping the vision encoder frozen. Additionally, it fused global and patch-averaged embeddings, enabling patch-level similarity computation with text and supporting open-vocabulary semantic segmentation. UniCLIP [20] introduced a multi-positive NCE (MP-NCE) loss, which independently evaluates multiple positive pairs per image. Building on these advances, our approach integrates a frozen, fine-grained vision encoder with trainable Transformer layers, following LiT and dino.txt. We also adopt MP-NCE loss to align images with multiple text representations effectively.
2.2 Vision-language alignment in radiology
Since the introduction of contrastive learning in radiology [42], aligning CXR images with radiology reports has become an active research area. GLoRIA [12] focused on local alignment using cross-attention between word-level text embeddings and patch-level image features. MGCA [33] employed both report-level and token-level embeddings to extract multi-granular features, and BioViL-T [1] similarly relied on token-level embeddings. Nevertheless, segmenting reports into individual words or tokens often fails to capture their full semantic meaning.
Due to the complexity of medical image–report relationships, alignment interpretability is essential for clinical use and is commonly addressed using attention maps. MedKLIP [36] and KAD [41], for example, used RadGraph [14] to extract report features and employed attention maps for tasks like grounding and segmentation. In addition to VL alignment, G2D [21] G2D [21] aggregates attention maps in addition to VL alignment to generate pseudo masks, which are used as pixel-wise pretext supervisory signals during pre-training. CARZero [19] also used attention maps when leveraging cross-attention alignment for zero-shot tasks, incorporating LLM-based prompt alignment to standardize reports. Despite their utility, attention maps have limitations: they often highlight irrelevant regions due to softmax activation, but removing softmax is not ideal as raw logits are unnormalized and uncentered. Additionally, variation in the norms of query and key embeddings leads to inconsistent similarity values across different image-text pairs. An example and a detailed discussion of the limited explainability of attention maps are provided in Appendix A. In contrast, our approach enhances explainability with VL-CABS, aligning visual patches and text embeddings. The resulting maps offer intuitive and consistent measures of fine-grained image-text similarity.
3 Methods
3.1 Finding-sentence extraction
Radiology reports contain diverse types of information, including clinical history, observations, comparative analysis with prior studies, and diagnostic impressions. Encoding the entire report into a single text embedding often fails to capture this complexity. CARZero [19] addressed this by using an LLM to extract relevant sentences and introducing a prompt alignment strategy based on the template “There is [disease]” for consistency between training and inference. Similarly, we use an LLM to extract such sentences, which we refer to as finding-sentences. These are generated using a prompt that follows a predefined structure, such as “There is [finding] of [location],” and are segmented into minimal semantic units containing the finding name, presence (or uncertainty), and location. The full prompt is provided in Appendix D. Each image is paired with multiple finding-sentences during training, as illustrated in Figure 3 (a). For zero-shot inference, we apply prompt alignment by prepending “There is” to text descriptions of findings and anatomical regions.
Figure 3: The overall framework of RadZero. (a) Finding-sentence extraction using an LLM. (b) Computation of the similarity logit, $l_{ij}^{n}$, between image $I_{i}$ and finding-sentence $S_{j}^{n}$. W-sum and cos-sim denote weighted sum and cosine similarity, respectively. (c) Computation of MP-NCE loss ($\mathcal{L}_{I}$) and InfoNCE loss ($\mathcal{L}_{T}$) from the similarity logit matrix. (d) Zero-shot inference pipeline.
3.2 Vision-language alignment with similarity based cross-attention
3.2.1 Model architecture
To leverage the advantages of vision encoder pre-training, we adopt the approach of LiT [38] by freezing a pre-trained vision encoder in contrastive learning. In Vision Transformers [9] such as DINOv2 [26], interpolating the positional embeddings allows for increased input image resolution [30]. Building on this property, we train our model with high-resolution images. To embed the output of the vision encoder, we add $K$ trainable Transformer layers, as proposed by Jose et al. [18]. For the text encoder, we use a pre-trained Sentence-BERT [29], which is fine-tuned during training, to extract embeddings for each finding-sentence. The model architecture is illustrated in Figure 3 (b).
3.2.2 Vision-language cross-attention based on similarity
We propose VL-CABS, a cosine similarity-based cross-attention mechanism for computing similarity logits. By directly employing cosine similarity between the text and visual patch embeddings, we obtain VL similarity scores that are well-defined in range and centered at zero. This consistent scaling allows for fair comparisons across different image–text pairs and significantly enhances explainability through the visualization of VL similarity maps. It also enables single thresholding, offering new possibilities for open-vocabulary semantic segmentation.
The proposed method operates on a mini-batch of size $B$, pairing each image $I_{i}$ ($i=1,\dots,B$) with $N_{i}$ associated finding-sentences $\{S_{i}^{n}\}_{n=1}^{N_{i}}$. Each image is processed by a vision encoder $f_{v}$ followed by trainable layers $f_{a}$, yielding a sequence of embeddings $[\,v_{i0},\,v_{i1},\dots,v_{iL}\,]=f_{a}(f_{v}(I_{i})),$ where $v_{i0}\in\mathbb{R}^{D}$ corresponds to the [CLS] token and $\{v_{ik}\}_{k=1}^{L}\subset\mathbb{R}^{D}$ are patch embeddings, with $D$ denoting the embedding dimension and $L$ the total number of patches. Each finding-sentence $S_{j}^{n}$ is encoded into a sentence-level embedding $t_{j}^{n}=f_{t}(S_{j}^{n})\in\mathbb{R}^{D}$ using a text encoder $f_{t}$.
To compute VL similarity, we ℓ2-normalize all embeddings as $\bar{v}_{ik}=v_{ik}/\|v_{ik}\|_{2}$ and $\bar{t}_{j}^{n}=t_{j}^{n}/\|t_{j}^{n}\|_{2}$, and calculate the scaled cosine similarity between each patch and sentence as $s_{ijk}^{n}=(\bar{v}_{ik})^{\top}\bar{t}_{j}^{n}\cdot\exp(\tau)\quad(k=0,\dots,L),$ where $\tau$ is a learnable temperature. These scores are converted into attention weights using softmax over the patch index $k$: $a_{ijk}^{n}=\exp(s_{ijk}^{n})/{\sum_{m=0}^{L}\exp(s_{ijm}^{n})}.$ A sentence-specific attended vision embedding is computed as the weighted sum $v_{ij}^{n}=\sum_{k=0}^{L}a_{ijk}^{n}v_{ik}$, which is then ℓ2-normalized as $\bar{v}_{ij}^{n}=v_{ij}^{n}/\|v_{ij}^{n}\|_{2}$. The global similarity logit between image $I_{i}$ and sentence $S_{j}^{n}$ is given by $l_{ij}^{n}=({\bar{v}_{ij}^{n}})^{\top}\bar{t}_{j}^{n}\cdot\exp(\tau).$ The corresponding patch-level similarity map is $M_{ij}^{n}=[\,s_{ij1}^{n},\dots,s_{ijL}^{n}\,]$.
3.3 Multi-positive contrastive learning
Although CARZero also uses prompt templates for training, it suffers from instability due to randomly selecting one sentence for each image at every training step. To utilize all $N$ finding-sentences matched to each image at every step, we adopt multi-positive NCE (MP-NCE) loss [20] which treats positive pairs independently in order to amplify the loss contributions from each positive pair. A visualization of our contrastive loss can be seen in Figure 3 (c). Let $N_{T}=\sum_{i=1}^{B}N_{i}$ be the total number of finding-sentences in a mini-batch. For the $i$-th image, the number of positive and negative finding-sentences are $N_{i}$ and $N_{T}-N_{i}$, respectively. The MP-NCE loss can be computed as follows:
$\mathcal{L}_{I}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{\exp(l_{ii}^{n})}{\exp(l_{ii}^{n})+\sum\limits_{j\neq i}^{B}\sum\limits_{m=1}^{N_{j}}\exp(l_{ij}^{m})}$ (1)
For each finding-sentence $S_{i}^{n}$, there is one positive image $I_{i}$ and $B-1$ negative images. The corresponding InfoNCE loss [25] is computed as follows:
$\mathcal{L}_{T}=-\frac{1}{N_{T}}\sum_{i=1}^{B}\sum_{n=1}^{N_{i}}\log\frac{\exp(l_{ii}^{n})}{\exp(l_{ii}^{n})+\sum\limits_{j\neq i}^{B}\exp(l_{j,i}^{n})}$ (2)
The final objective function is the sum of $\mathcal{L}_{I}$ and $\mathcal{L}_{T}$: $\mathcal{L}=\mathcal{L}_{I}+\mathcal{L}_{T}.$
3.4 Zero-shot inference
The similarity logit between an image $I_{i}$ and a sentence $S_{j}^{n}$, denoted by $l_{ij}^{n}$, is converted into a similarity probability $\hat{l}_{ij}^{n}=\sigma(l_{ij}^{n})$ via a sigmoid function, and used for zero-shot classification.
For grounding and segmentation, we reshape the patch-level similarity map $M_{ij}^{n}=[s_{ij1}^{n},\dots,s_{ijL}^{n}]$ into a $\sqrt{L}\times\sqrt{L}$ square map, and resize it to the original image resolution via bilinear interpolation. To account for preprocessing such as padding and resizing, this interpolation is applied accordingly. A final element-wise sigmoid activation is applied to obtain the pixel-level similarity map $\hat{M}_{ij}^{n}=\sigma(\text{bilinear}(M_{ij}^{n}))$, which we refer to as the VL similarity map, used for zero-shot grounding and segmentation. The VL similarity map is derived from the cosine similarity between vision patches and text embeddings, and since we do not modify the embedding space beyond applying ℓ2-normalization and adjusting the temperature, it can be directly interpreted as the similarity between each image pixel and the text. The zero-shot inference process is illustrated in Figure 3 (d).
4 Experiments
4.1 Training dataset
MIMIC-CXR [16]
We train our model using the MIMIC-CXR dataset for VL alignment. MIMIC-CXR comprises 377,110 CXR images from 227,835 radiographic studies involving 65,379 patients. Each study includes a radiology report and one or more CXR images in either frontal or lateral views. Images are sourced from MIMIC-CXR-JPG [17], and only the findings and impression sections of reports are extracted using the official codebase [^1] https://github.com/MIT-LCP/mimic-cxr. All view positions are considered, and the official dataset split is followed. As described in Section 3.1, finding-sentence extraction is applied, with each study containing an average of 6.45 such sentences. Studies without extracted finding-sentences are discarded, resulting in 352,875 training images and 2,852 for validation.
4.2 Test datasets
Open-I (OI) [7] contains 3,851 radiology reports and 7,470 CXR images with multi-label annotations for 18 diseases. PadChest (PC) [4] comprises 160,868 CXR images from 67,000 patients, with 192 labels showing a long-tailed distribution. Following [19], we use 39,053 samples annotated by board-certified radiologists. Additionally, PadChest20 (PC20), introduced in [19], serves as a test set for rare disease evaluation, consisting of 20 classes with fewer than 10 samples each. ChestXray14 (CXR14) [34] provides official test set with 22,433 images labeled for 14 diseases. CheXpert (CXP) [13] includes a test set of 500 patients’ images annotated by five board-certified radiologists. Following [19], we evaluate classification on five observations: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. ChestXDet10 (CXD10) [22], a subset of CXR14, contains 542 images with bounding box annotations for 10 diseases in the official test set. SIIM [37] pneumothorax dataset provides segmentation masks for 11,582 CXRs; we adopt the test split from [33], which includes 1,704 images with 458 positives. RSNA [31] pneumonia dataset consists of 29,700 frontal CXRs with bounding box annotations; we use the test set from [36] containing 5,337 images, including 1,218 positives. MS-CXR [2] consists of 1,153 image-phrase-bounding box triplets, with images sourced from MIMIC-CXR. The bounding boxes annotated to specific phrases in the report enable more detailed grounding, referred to as phrase grounding. For fair evaluation on the 167-image test set released by [6], where each phrase maps to a single bounding box, we exclude these images from the training set described in Sec. 4.1.
4.3 Evaluation metrics
AUC, or area under the ROC curve, is adopted to evaluate zero-shot classification on multi-label test datasets. Pointing game [40], which determines whether the coordinates of the maximum value falls within the corresponding bounding box, is employed as the grounding metric. Dice score serves as a standard evaluation metric for segmentation. Following [36], we compute the Dice score using only positive samples and optimize the segmentation threshold to maximize the score. Threshold search intervals are 0.01 for sigmoid and 0.001 for softmax, depending on the feature map’s activation function. Pixel-wise AUC (Pix-AUC) computes AUC at pixel-level to evaluate the quality of the segmentation probability map. To account for both sensitivity and specificity in mask prediction, we incorporate both positive and negative samples. For fine-grained tasks such as grounding and segmentation, predictions are interpolated back to the original image size before evaluation.
4.4 Implementation details
We adopt XrayDINOv2 [5] as the pre-trained vision encoder, which is trained in a unimodal setting using CXR images based on DINOv2 [26]. While the vision encoder was trained with an image resolution of 224, we increase it to 518 for our experiments. The patch size of $14\times 14$ leads to $37\times 37$ patches, yielding a vision patch length $L$ of 1369. The text encoder is MPNet ("all-mpnet-base-v2") [29], initialized with pre-trained parameters. The trainable Transformer layers consist of two randomly initialized layers ($K=2$), with a hidden dimension of 768, matching the embedding size of both the vision and text encoders. While the vision encoder remains frozen, all other parameters are trainable. Following [28], the learnable temperature parameter $\tau$ is initialized to $\log(1/0.07)$. The details of model training can be found in Appendix B. The LLM used for extracting finding-sentences is "Llama-3.3-70B-Instruct" [10], deployed in a private computing environment.
5 Results
5.1 Zero-shot evaluation
  Method Open-I (OI) PadChest (PC) PadChest20 (PC20) ChestXray14 (CXR14) CheXpert (CXP) ChestXDet10 (CXD10) SIIM RSNA
GLoRIA [12] 0.589 0.565 0.558 0.610 0.750 0.645 - -
BioViL-T [1] 0.702 0.655 0.608 0.729 0.789 0.708 - -
MedKLIP [36] 0.759 0.629 0.688 0.726 0.879 0.713 0.897 0.869
KAD [41] 0.807 0.750 0.735 0.789 0.905 0.735 - -
CARZero [19] 0.838 0.810 0.837 0.811 0.923 0.796 0.924 0.747
RadZero (224px) 0.851 0.841 0.879 0.807 0.903 0.785 0.914 0.839
RadZero 0.847 0.841 0.871 0.804 0.900 0.787 0.924 0.834
Table 1: Zero-shot classification AUROC scores on public CXR datasets. For fair comparison, we also report the results of low-resolution (224×224) version of RadZero.
Classification.
Table 1 compares RadZero with SOTA models on public test datasets. For the five datasets evaluated in CARZero [19], we report their published results. For SIIM and RSNA, we independently evaluated two open-source models. RadZero achieved new SOTA performance on OI and PC, irrespective of image resolution. In the long-tailed PC dataset with 192 classes, RadZero outperformed CARZero by 3.1 percentage points, demonstrating strong generalization in zero-shot classification. Notable gains are also observed in PC20, which focuses on rare diseases, suggesting that VL-CABS is particularly effective for infrequent conditions. On datasets where RadZero failed to rank first, MedKLIP performed best on RSNA, while CARZero led on CXR14, CXP, and CXD10. However, MedKLIP underperformed on the latter datasets, and CARZero underperformed on RSNA. In contrast, RadZero showed results comparable to the top-performing models across all datasets. Interestingly, the lower-resolution RadZero (224px) even outperformed RadZero: potentially due to the pre-trained vision encoder. as discussed in Sec. 5.4.
The representative classification metric shown in Figure 2 is the average AUC across all datasets. RadZero established a new SOTA, outperforming CARZero by 1.5 percentage points, a gain attributable to our training strategy that incorporates multi-positive contrastive learning to enhance the diversity of both positive and negative samples per image.
Grounding.
  Method Mean ATE CALC CONS EFF EMPH FIB FX MASS NOD PTX
GLoRIA [12] 0.367 0.479 0.053 0.737 0.528 0.667 0.366 0.013 0.533 0.156 0.143
KAD [41] 0.391 0.646 0.132 0.699 0.618 0.644 0.244 0.199 0.267 0.316 0.143
BioViL-T [1] 0.351 0.438 0.000 0.630 0.504 0.846 0.390 0.026 0.500 0.000 0.171
MedKLIP [36] 0.481 0.625 0.132 0.837 0.675 0.734 0.305 0.224 0.733 0.312 0.229
CARZero [19] 0.543 0.604 0.184 0.824 0.782 0.846 0.561 0.184 0.700 0.286 0.457
RadZero (224px) 0.537 0.604 0.211 0.806 0.813 0.795 0.451 0.197 0.767 0.325 0.400
RadZero 0.622 0.646 0.368 0.824 0.857 0.872 0.585 0.250 0.767 0.506 0.543
Table 2: Zero-shot grounding results (pointing game accuracy) on CXD10. Lesion abbreviations can be found in the Appendix E.
  Method MS-CXR
BioViL-T [1] 0.719
MedKLIP [36] 0.407
CARZero [19] 0.749
RadZero (224px) 0.832
RadZero 0.844
Table 3: Zero-shot phrase grounding results (pointing game accuracy) on MS-CXR.
Table 2 presents zero-shot grounding results on CXD10. We adopted the pointing game scores reported by CARZero for all models except BioViL-T, which we evaluated using its released weights. RadZero achieved the highest average score across all diseases, outperforming CARZero by 0.079. Per-lesion analysis showed that RadZero achieved the best performance in all classes except consolidation, indicating that the proposed VL-CABS effectively captures local alignment between text and image patches regardless of disease type. Furthermore, RadZero efficiently supports higher input resolutions, enabling more precise localization.
Table 3 reports zero-shot phrase grounding reports, with pointing game accuracy as the evaluation metric. We evaluated all baselines using publicly available models. RadZero achieved the highest score of 0.844, demonstrating accurate interpretation of text phrases. The strong performance of RadZero (224px) suggests that the gains are largely attributable to the effectiveness of VL-CABS rather than the higher input resolution.
Segmentation.
  Method RSNA SIIM
Dice Dice Pix-AUC
GLoRIA [12] 0.347* - -
BioViL [3] 0.439* - -
MedKLIP [36] 0.465* 0.044 0.648
G2D [21] 0.477 0.051 -
CARZero [19] 0.540 0.100 0.856
CARZero (logits) 0.529 0.081 0.928
RadZero (224px) 0.562 0.121 0.943
RadZero 0.546 0.171 0.947
  MGCA [33] (1%) 0.513 0.144 0.752
MGCA (10%) 0.571 0.238 0.856
MGCA (100%) 0.578 0.305 0.976
Table 4: Zero-shot segmentation results. Values with are from [36] and from [21]
Table 4 summarizes zero-shot segmentation results on SIIM and RSNA. To benchmark against supervised models, we fine-tuned MGCA with varying proportions of training data; percentages in parentheses indicate the amount used.
Among zero-shot models, RadZero achieved the highest Dice scores on both datasets. On SIIM, it outperformed CARZero by 71% demonstrating superior segmentation capability. The smaller margin on RSNA is likely due to its coarser annotations—bounding boxes rather than pixel-level masks—which limit the advantages of RadZero’s fine-grained VL similarity maps. As in phrase grounding, the results of RadZero (224px) suggest that the superior performance is not merely driven by higher resolution.
RadZero also remained competitive against fine-tuned models. It outperformed MGCA (1%) on both datasets, demonstrating the effectiveness of zero-shot segmentation. Although it did not surpass MGCA (10%) or (100%), RadZero requires no mask labels, enabling broader generalization beyond fixed vocabularies, as further discussed in Sec. 5.3.
SIIM’s detailed annotations also allow for Pix-AUC evaluation. RadZero achieved the highest score, exceeding even MGCA (10%), indicating well-calibrated VL similarity maps that distinguish positive from negative regions. In contrast, MedKLIP and CARZero, both relying on attention maps, performed worse with scores of 0.648 and 0.856, respectively. For a fair comparison, we also evaluated CARZero’s pre-softmax logits (CARZero (logits)), which improved performance but still fell short of RadZero. Notably, CARZero (logits) underperformed its own Dice score (0.081 vs. 0.100) despite threshold tuning, as expected from the inconsistent scaling of dot-product similarity. In contrast, VL-CABS directly encode pixel-level text–image similarity, allowing low values for negative samples. This contributed to its superior Pix-AUC, as further supported in Sec. 5.2.
Figure 4: VL similarity maps of CXR images from CXD10, representing (a) normal, (b) fibrosis, and (c) effusion in the right lung. The value at the top-right corner represent the similarity probability $\hat{l}$ between each CXR image and the text prompt (bottom-right corner).
Figure 5: Open-vocabulary semantic segmentation results: (a), (b) for findings and (c) for anatomical regions. The CXR images and bounding box labels are from CXD10. The segmentation thresholds were set to 0.7 for (a) and (b), and 0.4 for (c).
5.2 VL similarity map analysis
Figure 5 illustrates that RadZero effectively aligned visual and textual representations with VL-CABS. The outputs, VL similarity map $\hat{M}$ and probability $\hat{l}$, offer both interpretable visualizations and quantitative metrics. For the normal image Figure 5 (a), the model assigned low similarity (0.072) to the prompt “There is atelectasis,” with a dark VL similarity map, indicating weak alignment between vision tokens and the text embedding, In contrast, “There is no atelectasis” (0.921) and “The lungs are clear” (0.846) yielded bright activations across lung fields, reflecting strong alignment.
Figure 5 (b) shows fibrosis, and “There is fibrosis” resulted in high similarity (0.970) with strong activations in the affected lung. Prompts indicating normality received much lower scores (0.093 and 0.014) and darker VL similarity maps, clearly distinguishing abnormal from normal descriptions.
Figure 5 (c) highlights RadZero’s ability to distinguish anatomical descriptions. For right-sided pleural effusion, the model assigned high similarity (0.907) to “There is effusion,” with bright activations in the correct region. Notably, “There is right effusion” (0.946) scored even higher, indicating accurate localization, while “There is left effusion” scored much lower (0.125) and a dark VL similarity map, showing that the model correctly distinguishes between left and right lung regions.
Overall, these results underscore the explainability of VL-CABS. The similarity probability is verifiable at the pixel level, enabling spatially grounded explanations. By explicitly revealing how conclusions are derived, RadZero offers enhanced interpretability in the context of disease diagnosis.
5.3 Open-vocabulary semantic segmentation
Figure 5 presents open-vocabulary semantic segmentation results for both findings and anatomical regions. Segmentation masks were generated by thresholding the VL similarity map $\hat{M}$ for each text prompt; in cases of overlapping predictions, the prompt with the highest similarity was assigned. In Figure 5 (a), RadZero successfully localized lesions based on text queries, though some segmentation masks extended beyond ground truth boxes, indicating room for improvement. Notably, certain incorrect predictions captured clinically relevant features that were not explicitly annotated: in Figure 5 (b), a chest tube was reasonably associated with “pneumothorax.” Figure 5 (c) further demonstrates RadZero’s ability to segment anatomical structures without supervision, inferring approximate spatial regions from text despite imprecise boundaries. These results highlight the potential of VL-CABS for zero-shot open-vocabulary semantic segmentation and RadZero’s capacity to align textual descriptions with medical imagery. Additional qualitative examples are provided in Appendix F.
5.4 Ablation Studies
Method Similarity Trainable layers MP Res. Classification Grounding Segmentation
OI PC PC20 CXR14 CXP CXD10 CXD10 MS-CXR SIIM
(a) dot-product Linear 224 0.839 0.824 0.853 0.805 0.896 0.792 0.472 0.784 0.078
(b) $\cos$ Linear 224 0.843 0.830 0.863 0.805 0.902 0.786 0.483 0.790 0.078
(c) $\cos$ 2 Transformer 224 0.845 0.832 0.860 0.808 0.895 0.793 0.539 0.838 0.099
(d) RadZero (224px) $\cos$ 2 Transformer 224 0.851 0.841 0.879 0.807 0.903 0.785 0.537 0.832 0.121
RadZero $\cos$ 2 Transformer 518 0.847 0.841 0.871 0.804 0.900 0.787 0.622 0.844 0.171
LiT [38] - Linear 224 0.768 0.769 0.775 0.764 0.854 0.735 - - -
dino.txt [18] - 2 Transformer 224 0.834 0.816 0.837 0.797 0.901 0.770 0.121 0.174 0.021
CARZero [19] - Transformer Dec. 224 0.827 0.815 0.877 0.795 0.889 0.770 0.437 0.743 0.072
Table 5: Ablation study of model architecture components. “MP” denotes multi-positive.
Ablation study on RadZero components.
1) Similarity function: Comparing (a) and (b) in Table 5 shows the effect of using cosine similarity instead of scaled dot-product [32] for VL alignment. Cosine similarity, which better aligns with inference-time VL similarity maps, improves classification and slightly enhances grounding. 2) Trainable parameters: (b) and (c) compare a linear layer and a two-layer Transformer, with the image encoder frozen. Transformer layers yielded consistent gains across all tasks, particularly in grounding (0.483 → 0.539) and segmentation (0.078 → 0.099). 3) Multi-positive pairs: The difference between (c) and (d) lay in the use of multi-positive contrastive pairs. (d) improved classification (e.g., PC20: 0.860 → 0.879) and segmentation (0.099 → 0.121), highlighting the advantage of richer supervision. 4) Image resolution: RadZero and (d) shared the same architecture, except that RadZero used higher resolution inputs (518 vs. 224). This change substantially improved grounding (0.537 → 0.622) and segmentation (0.121 → 0.171), showing the importance of high-resolution features for spatially localized tasks. In classification, (d) outperformed RadZero: likely due to the vision encoder (XrayDINOv2[5]) being pre-trained at 224 pixels, indicating that using an encoder pre-trained at higher resolutions may further enhance the performance.
Comparison among different VL alignment approaches.
Table 5 compares RadZero with alternative VL alignment methods, keeping all settings identical except for VL feature fusion and loss computation. LiT, which uses a [CLS] embedding for alignment, showed limited classification performance and was incapable of grounding or segmentation. dino.txt improved classification through additional Transformer layers, but its mean pooling constrained grounding and segmentation performance. CARZero introduced a cross-attention decoder, enhancing performance on those tasks. However, when compared to the RadZero ablations, (b) outperformed CARZero across most metrics, showing that VL-CABS alone is sufficiently effective.
6 Conclusion
In this work, we introduced RadZero, a novel VL alignment model for radiology that achieved strong zero-shot performance in classification, grounding, and segmentation. Central to RadZero is VL-CABS, which computes image-text similarity at the patch-level to improve interpretability. Combined with multi-positive contrastive training, VL-CABS enabled effective representation learning without pixel-level annotations, and the support for high-resolution inputs further boosted performance. Extensive evaluations on public chest radiograph benchmarks showed that RadZero outperformed SOTA methods. VL similarity map analysis highlighted the enhanced explainability of VL-CABS by providing transparent rationales for how conclusions are derived. Qualitative assessments further demonstrated RadZero’s potential for open-vocabulary semantic segmentation.
Despite its impressive results, RadZero has limitations that indicate areas for future research. The performance lag on certain datasets highlights the need for improved generalization. Its reliance on the pre-trained vision encoder is also a constraint. Fine-tuning with class labels could facilitate the development of a high-performance explainable classifier. Extending VL-CABS to general image domains offers potential for more interpretable VLMs across diverse visual tasks.
Appendix A Are Attention Maps of Medical VLMs Explainable?
Explainability is critical for clinical deployment of deep learning models, and medical VLMs have adopted attention maps as their de facto interpretable features [36, 19, 41]. However, attention maps alone reveal only where the model is focusing, not why it makes a particular prediction. Without the underlying image–text similarity scores, such heatmaps lack interpretability. Interpreting attention maps for complex text queries can be unintuitive and may depend on access to ground-truth labels.
Figure 6: Attention maps from CARZero [19] Image–text similarity probabilities are obtained by applying a sigmoid function to the classification logits.
To illustrate the challenges of interpreting attention maps, Figure 6 visualizes the attention outputs of CARZero [19] for different image–text pairs, with the overlaid values indicating the corresponding similarity probabilities. In (a), the prompt “There is right effusion” produces a heatmap focused on the right lower lung and a high similarity probability (0.560), intuitively linking the attended region to the predicted finding. In (b), when queried with “There is left effusion” on the same image, the attention map still highlights the right lower lung. While we can observe where the model attends, the attention map alone offers no clear explanation for why that region is relevant. The image–text similarity probability for the prompt “There is left effusion” is low (0.048), which—when considered alongside the attention map—can be interpreted as indicating that the attended right lung region does not support the presence of left-sided effusion. However, how should we interpret the fact that the model attends to the right lung when queried about left effusion? Making sense of this behavior requires access to the ground truth: effusion is present in the right lower lung but absent on the left. With this context, one might infer that the model correctly identifies effusion on the right, recognizes that it does not match the left-sided query, and therefore assigns a low similarity score.
This example underscores the need to interpret attention maps in conjunction with image–text similarity scores. For complex query texts, meaningful interpretation often depends on knowing the ground truth—a significant limitation in medical VLMs where such labels are frequently unavailable. This reliance undermines the standalone explainability of attention visualizations. To address these limitations, we propose Vision–Language Cross-Attention Based on Similarity (VL-CABS), a framework that enables transparent inspection of the vision–language decision process. Detailed examples and analysis of its application in RadZero are provided in Section 5.2.
Appendix B Model Training Details
RadZero is trained for 20 epochs with an early stopping patience of 5 epochs, selecting the best model based on validation loss. We employ the AdamW optimizer with a learning rate of 0.0001, following a cosine decay scheduler, with 50 warm-up steps, a weight decay of 0.05, and gradient clipping set to 1.0. Training is conducted with a global batch size of 256 using distributed data parallel (DDP) on four H100 GPUs for 16 hours.
Appendix C Ablation Studies
We conducted ablation studies to assess the impact of key design choices by selectively modifying parts of our approach. The zero-shot tasks for evaluation included classification (class.), grounding (ground.), phrase grounding (phrase.), and segmentation (seg.). Classification was tested on the PadChest dataset, known for its highly imbalanced (long-tailed) label distribution, with AUC as the evaluation metric. Grounding and phrase grounding were evaluated using the pointing game on the ChestXDet10 and MS-CXR test sets, respectively. Segmentation performance was measured by the Dice score on the SIIM dataset. For the ablation study, the default batch size and maximum number of epochs were set to 128 and 10, respectively.
Impact of view position.
Table 10 shows the performance variation based on the view position of CXR images. We compared two models: one trained exclusively on frontal view images from MIMIC-CXR and another trained on both frontal and lateral views. The model trained on all view positions consistently outperformed the frontal-only model, suggesting that it effectively learned to interpret lateral images, enhancing overall robustness.
Impact of image encoder and resolution.
Table 10 presents the impact of the image encoder and image resolution on model performance. We compared M3AE [11], RadDINO [27], and XrayDINOv2 [5] using image resolutions of 224 and 518. For M3AE, we used the pre-trained model released by Lai et al. [19], and all encoders were frozen during training.
DINOv2-based encoders significantly outperformed M3AE across all tasks, demonstrating the effectiveness of the DINOv2 pre-training strategy for chest X-ray representation learning. Comparing XrayDINOv2 at resolutions of 224 and 518, we observe that higher image resolution improves fine-grained tasks such as grounding and segmentation. RadDINO and XrayDINOv2 showed similar performance, suggesting that our approach is effectively applied to models trained with the DINOv2 strategy on chest X-ray images.
Impact of trainable vision layer architecture.
  View Position class. ground. phrase. seg.
Frontal 0.831 0.604 0.838 0.161
All View 0.841 0.622 0.844 0.171
Table 6: Impact of view position.
  Image Image class. ground. phrase. seg.
Encoder Resolution
M3AE [11] 224 0.703 0.284 0.647 0.027
RadDINO [27] 518 0.850 0.610 0.844 0.144
XrayDINOv2 [5] 224 0.841 0.548 0.832 0.118
XrayDINOv2 518 0.841 0.622 0.844 0.171
Table 7: Impact of image encoder and resolution.
  Model class. ground. phrase. seg.
Linear 0.826 0.549 0.826 0.100
1 Transformer layer 0.835 0.585 0.832 0.158
2 Transformer layers 0.841 0.622 0.844 0.171
Table 8: Impact of trainable vision layer architecture.
  Image class. ground. phrase. seg.
Encoder
BioBERT 0.842 0.582 0.832 0.127
MPNet 0.841 0.622 0.844 0.171
Table 9: Impact of text encoder.
  Batch size class. ground. phrase. seg.
64 0.835 0.583 0.826 0.165
128 0.840 0.594 0.850 0.177
256 0.841 0.622 0.844 0.171
Table 10: Impact of batch size.
Table 10 presents the impact of different trainable layers in the image encoder. The commonly used linear layer showed relatively lower performance across tasks. In contrast, two transformer layers achieved the best results in classification, grounding, and phrase grounding.
Based on this observation, RadZero was designed with two transformer layers added to the vision encoder. This improvement is likely due to the transformer’s ability to attend to all patch embeddings, capturing richer semantic information.
Impact of text encoder.
Table 10 presents the performance of different text encoders used during training. We compared MPNet and BioBERT, where BioBERT was fine-tuned on clinical reports by CARZero [19].
While MPNet showed slightly lower performance in classification, it achieved notable improvements in phrase grounding and segmentation, demonstrating its effectiveness in tasks requiring fine-grained text-image alignment.
Impact of batch size.
Table 10 presents the impact of batch size on model performance during training. To ensure a fair comparison, we maintained a consistent total number of training steps by adjusting the number of epochs: 5 for a batch size of 64, 10 for 128, and 20 for 256.
We observed that a batch size of 64 resulted in lower performance across all tasks. While the model trained with a batch size of 128 performed reasonably well, its zero-shot grounding performance was notably lower than that of the 256 batch size model. As a result, we selected 256 as the final batch size.
This trend aligns with the well-known impact of batch size in contrastive learning, where larger batch sizes generally improve representation learning by providing more diverse negative samples, leading to better alignment and discrimination.
Appendix D Prompt for finding-sentence extraction.
As shown in Figure 7, the prompt instructs the LLM to extract clinically relevant minimal semantic units in the form of sentences from radiology reports. Finding-sentences are standardized through prompt alignment to follow a “There is” format, with a one-shot example enhancing extraction accuracy and guiding the model to identify both findings and their corresponding anatomical locations in a structured manner.
  Abbreviation Description
ATE Atelectasis
CALC Calcification
CONS Consolidation
EFF Effusion
EMPH Emphysema
FIB Fibrosis
FX Fracture
MASS Mass
NOD Nodule
PTX Pneumothorax
UL Upper Lobe
ML Mid Lobe
LL Lower Lobe
Rclav Right Clavicle
Lclav Left Clavicle
RULZ Right Upper Lung Zone
RMLZ Right Mid Lung Zone
RLLZ Right Lower Lung Zone
LULZ Left Upper Lung Zone
LMLZ Left Mid Lung Zone
LLLZ Left Lower Lung Zone
RCPA Right Costophrenic Angle
LCPA Left Costophrenic Angle
HD Hemidiaphragm
RHD Right Hemidiaphragm
LHD Left Hemidiaphragm
Table 11: Abbreviations for lesions and anatomical regions.
Figure 7: Prompt design for extracting finding-sentences with LLM.
Appendix E Abbreviation
Table 11 lists the abbreviations used for lesions and anatomical regions. The left column shows the abbreviated terms, and the right column gives their description.
Figure 8: Open-vocabulary semantic segmentation for anatomical regions. The CXR images are sourced from Open-I. The segmentation threshold was set to 0.4.
Appendix F Additional Visualization Results
Figure 8 and Figure 10 present open-vocabulary semantic segmentation results, following Sec. 5.3. These figures include images from Figure 1 for qualitative analysis, along with additional examples not shown in the introduction.
Figure 8 depicts segmentation of anatomical regions, which, while not perfect, generally align with appropriate locations. Figure 10 presents examples demonstrating RadZero’s potential for open-vocabulary semantic segmentation, including additional lesion types such as mass, fibrosis, and calcification.
Figure 10 presents VL similarity maps for 10 different findings, following the pipeline in Sec. 5.2. The highest similarity regions align well with the bounding boxes, even for multiple or small lesions. The similarity probability was above 0.5 for all findings except calcification. While the model correctly localized calcifications, the activated regions appeared as small bright spots, leading to a lower similarity probability of 0.45 due to the weighted sum calculation. This highlights a limitation of RadZero, suggesting the need for further refinement in future work.
Figure 9: Open-vocabulary semantic segmentation for findings. The CXR images and bounding box labels are sourced from ChestXDet10. The segmentation threshold was set to 0.7.
Figure 10: VL similarity maps for 10 findings. Visualization of similarity maps generated by RadZero on the ChestXDet10 dataset. Red boxes indicate ground truth bounding boxes. The similarity probability $\hat{l}$ is shown in the top-right corner of each map.